OpenStudy (anonymous):

What is the discontinuity and zero of the function f(x) = the quantity of 3 x squared plus x minus 4, all over x minus 1?

OpenStudy (anonymous):

Discontinuity at (-1, 1), zero at (four thirds, 0) Discontinuity at (-1, 1), zero at (negative four thirds , 0) Discontinuity at (1, 7), zero at (four thirds , 0) Discontinuity at (1, 7), zero at (negtive four thirds, 0)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (freckles):

\[f(x)=\frac{3x^2+x-4}{x-1}\]

OpenStudy (anonymous):

Yes that's what it looks like.

OpenStudy (freckles):

To find when f=0 find when top=0 To find when f is discontinuous from when bottom=0

OpenStudy (anonymous):

What?

OpenStudy (freckles):

To find f=0 is to find the zeros. To find the zeros, find when top=0.

OpenStudy (freckles):

The top is 3x^2+x-4

OpenStudy (anonymous):

So set it equal to zero?

OpenStudy (freckles):

top=0 yes

OpenStudy (freckles):

and then you need to also find when bottom=0

OpenStudy (anonymous):

3x^2+x-4=0 and x-1=0

OpenStudy (freckles):

3x^2+x-4=0 will give you the zeros of f (aka when f=0) x-1=0 will give you the discontinuities Be careful though. If you get a zero of the top that is also a discontinuity of the bottom, then as of whole it is a discontinuity of the whole function, f(x).

OpenStudy (freckles):

Anyways what do you get when you solve 3x^2+x-4=0

OpenStudy (anonymous):

(x+1)(3x-4)=0

OpenStudy (freckles):

ok whick gives you x=? or x=?

OpenStudy (anonymous):

x=-1 and x=4/-3

OpenStudy (freckles):

Wait how did you get (x+1)(3x-4)?

OpenStudy (freckles):

I think you got the + and - mixed up

OpenStudy (anonymous):

so its (x-1) and (3x+4)?

OpenStudy (freckles):

3x^2+x-4 3x^2+4x-3x-4 x(3x+4)-1(3x+4) (3x+4)(x-1) yes.

OpenStudy (freckles):

so what do you get for x when you have (3x+4)(x-1)=0?

OpenStudy (anonymous):

x=1 and x=-4/3

OpenStudy (freckles):

But can x=1 be a zero?

OpenStudy (freckles):

of f?

OpenStudy (anonymous):

No

OpenStudy (freckles):

because f is actually discontinuous at x=1 because the bottom x-1 will be zero when x=1

OpenStudy (anonymous):

right so the answer is B?

OpenStudy (freckles):

no

OpenStudy (anonymous):

Oh

OpenStudy (freckles):

you said the function was discontinuous at x=1 and zero at x=-4/3 right?

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

only one choice says both of those things and it isn't B

OpenStudy (anonymous):

Its D

OpenStudy (freckles):

remember ordered pairs have the x first and y second

OpenStudy (freckles):

yes do you how to figure out the y of the discontinuity part?

OpenStudy (freckles):

we are done but I was just curious if you did

OpenStudy (anonymous):

No but I haven't learned to do that yet.

OpenStudy (freckles):

\[f(x)=\frac{(3x+4)(x-1)}{x-1}\] This what we get when we factor right?

OpenStudy (freckles):

now if you cancel the x-1's out we have \[f(x)=3x+4 , x \neq 1 \]

OpenStudy (freckles):

but if you did plug in 1 you get 3(1)+4=3+4=7

OpenStudy (freckles):

So that is how they got (1,7) for the discontinuity point

OpenStudy (anonymous):

Ah okay thanks for the help!

OpenStudy (freckles):

np

OpenStudy (anonymous):

Can you check my other answers then?

OpenStudy (freckles):

i could probably check a few

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

What is the discontinuity of the function f(x) = the quantity of x squared minus 4 x minus 12, all over x plus 2? (-6, 0) (6, 0) x (-2, -8) (2, -4)

OpenStudy (freckles):

this question can be answered by you if you understand what i did what do you think we need to do first?

OpenStudy (freckles):

obviously it is discontinuous where?

OpenStudy (freckles):

oh that x is suppose to symbolize your choice?

OpenStudy (anonymous):

The first thing I did was set the top part equal to zero and factored then I set the bottom part equal to zero.

OpenStudy (freckles):

\[\frac{(x-6)(x+2)}{x+2}\] ?

OpenStudy (anonymous):

yes thats what I got

OpenStudy (freckles):

and since the bottom equals 0 when x=-2 and when you cancel the (x+2)'s from the top and bottom you have f(x)=x-6 where x does not equal -2 but if you do plug in -2 you get?

OpenStudy (anonymous):

-8

OpenStudy (freckles):

 right if you wanted to know what this graph looks like notice that it is just the line y=x-6 with a hole at the point (-2,-8) since the function does not exist there